Operations Report – January 11th

Crew 202 Operations Report 11-JAN-2019

SOL: 13

Name of person filing report: Kasey Hilton

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator (hours run): 17hr 20min; Turned on last night (10Jan2019) at 16:05; Turned off this morning (11Jan2019) at 9:25; Turned on tonight (11Jan2019) at 17:15

Solar SOC – Turned on (10Jan2019) 84%; Turned off (11Jan2019) 100%; Turned on (11Jan2019) 81%

Diesel Reading – 65%

Propane Reading – 34%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – 40%; 220 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 55%; 165 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 20:41 to refill tank

Water Meter: 01399764

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: No

Hours: 67.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No, brakes are still not working correctly

Hours: 45.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 69.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brake fluid in Opportunity was replaced but the brakes are still not functioning properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: I am in the process of fixing the charging issues for suits #7 and #9; One of the electrical connections inside of suit #9 is loose and has to be soldered, I think the same thing is wrong with suit #7

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: The GreenHab door handle is loose but gets stuck when tightened; I am planning on checking the spring mechanism inside the handle

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report